It is given that 6 <= x <= 8 and 0.2 <= y <= 0.5. If z = x/y, find the limits in which z must lie.
Answer by jim_thompson5910(35256)

The largest x can be is 8
The smallest y can be is 0.2

So the largest z can be is 8/0.2 = 40

-------------------------------------------------------

The smallest x can be is 6
The largest y can be is 0.5

So the smallest z can be is 6/0.5 = 12

-------------------------------------------------------

If 6 <= x <= 8, 0.2 <= y <= 0.5, and z = x/y

Then 12 <= z <= 40